OpenStudy (anonymous):

derivative of y= (1/(e^(3x) +x^2))

OpenStudy (bahrom7893):

y=(e^(3x) +x^2)^(-1) y'=-(e^(3x)+x^2)^(-2)*[3e^(3x)+2x]

OpenStudy (anonymous):

could you help with f(z)=1/(e^z+1)^2

OpenStudy (anonymous):

and show steps?

OpenStudy (anonymous):

\[f(z)=\frac{1}{(e^z+1)^2}\]?

OpenStudy (anonymous):

you need chain rule for this. maybe easiest to rewrite in exponential notation as \[f(z)=(e^z+1)^{-2}\]

OpenStudy (anonymous):

then \[f'(z)=-2(e^z+1)^{-3}\times e^z\] by the chian rule

OpenStudy (anonymous):

clean up as \[f'(z)=\frac{-2e^z}{(e^z+1)^3}\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

could you help me with more of an application type?

OpenStudy (anonymous):

http://www.wiley.com/college/sc/hugheshallett/chap3.pdf Page 131 number 52 and 54